how do you do 2x+7y=-20 and y=3x+7

Answer by rfer(16322)

2x+7(3x+7)=-20
2x+21x+49=-20
23x=-69
x=-3
--------------
2*-3+7y=-20
-6+7y=-20
7y=-14
y=-14/7
y=-2
------------
y=3*-3+7
y=-9+7
y=-2